# Advancing Multimodal Data Fusion in Pain Recognition: A Strategy Leveraging Statistical Correlation and Human-Centered Perspectives

### v1

@@ IV-A Cross Validation
Analyzing the Leave-One-Out Cross-Validation (LOOCV) metrics presented in Table III, it is clear that both dual-modality and four-modality configurations outperform the single-modality training approach, thereby validating the benefits of data heterogeneity segmentation and human-centered modality segmentation. Specifically, the CNN-Attention model, when expanded to incorporate two modalities with statistical weighting, achieves a notable increase in accuracy (Acc.) from 0.800 to 0.907, and a lift in F1-score from 0.547 to 0.631. This improvement underscores the advantage of integrating multiple data sources over a singular modality framework.
@@
V Conclusion
This research venture has systematically unveiled the efficacy of integrating statistical methods and human-centered perspectives within the ambit of multimodal pain behaviours recognition, employing an array of deep learning architectures including convolutional neural networks (CNN), long short-term memory networks (LSTM), CNN-Attention networks and CNN with Multi-head Self Attention. The cornerstone of our exploration was to enhance the precision and utility of complex pain detection endeavors through the lens of statistical relevance and human-centered modality segmentation.
@@ V Conclusion
Furthermore, the research integrates statistical significance with a human-centered perspective, paving the way for explainable AI in pain recogintion. This approach delves into the distinct impact of each data modality on model outcomes, promoting a model of explainability that not only clarifies the workings of complex models but also enhances their adaptability and accuracy. This shift highlights our methodology’s broad applicability, not only in pain management but across diverse domains of human-centered computing, spotlighting its potential to revolutionize how we interact with and leverage AI technologies.
@@
This research intersects affective computing, multimodal data analysis, and pain assessment, aiming to enhance the precision and personalization of pain management through statistical methods. We conscientiously address the potential societal, environmental, and ethical impacts, focusing on aspects such as explainability, transparency, liability, fairness, efficacy, robustness, privacy, security, and sustainability.
@@ Acknowledgment
This work is partially supported by University College London Interaction Center, University College London, UK.

### v2

@@ IV-A Cross Validation
The Leave-One-Out Cross-Validation (LOOCV) metrics in Table III demonstrate that both dual-modality and four-modality configurations surpass the single-modality training approach, affirming the benefits of data heterogeneity segmentation and human-centered modality segmentation. Notably, the CNN-Attention model, when expanded to include two modalities with statistical weighting, shows a significant improvement in accuracy from 0.800 to 0.907 and an increase in F1-score from 0.547 to 0.631. This enhancement highlights the clear advantage of integrating multiple data sources over a single modality framework, indicating that the inclusion of diverse data types enriches the model’s performance in predicting outcomes.
@@
V Discussion
This research introduces an innovative approach to affective computing and multimodal data fusion by integrating statistical methods with human-centered computation. Our findings indicate that statistical algorithms can more effectively extract data representations across various modalities, which enhances our understanding and modeling of human behavior. Traditional machine learning methods often struggle with the complexity of human behavior and the interplay between different factors [24]. However, our approach successfully identifies key features linked to pain behavior by combining statistical correlations with human expert knowledge. This not only enhances model performance but also improves interpretability, providing valuable insights for the development of future interactive intelligent systems [25].
The study underscores the importance of data-driven statistical methods in pain recognition representation extraction, particularly in situations where complete reliance on patient self-reporting is not feasible. While self-reporting is a crucial means of pain recognition, it is limited by subjectivity, communication barriers, psychological factors, sociocultural influences, and feedback time delays[26]. Feature extraction based on statistical correlations provides an objective, continuous, and non-invasive method for pain representation extraction, complementing the limitations of self-reporting. This data-driven feature extraction can assist healthcare professionals in more accurately assessing patients’ pain conditions, especially when patients are unable to self-report[27].
Furthermore, given the substantial variability in human samples, personalisation has long been a potential barrier to digital healthcare. However, by utilising modality feature extraction driven by statistical correlations, our system can capture unique pain expression patterns for each individual, contributing to more precise and effective treatment strategies. This approach not only has the potential to improve patients’ quality of life but may also reduce the risk of drug abuse, particularly in the management of chronic pain[28].
Finally, the statistical correlation-driven multimodal fusion framework and its representation learning approach that we have demonstrated possess extensive application potential, with the possibility of further extension to other domains involving complex human-centred computing. This research provides a viable framework for integrating statistical methods, machine learning, and human expertise to address complex challenges of human-computer interaction. This interdisciplinary approach not only enhances technical performance but also augments its credibility and acceptability in practical applications. As technology continues to advance, we anticipate seeing more innovative applications based on this framework, ultimately realising more intelligent and humanised interactive systems. This integrated approach opens up new possibilities for future human-machine collaboration, with the potential to play a crucial role in various complex human-centred computing tasks.
VI Conclusion
This research venture has systematically unveiled the efficacy of integrating statistical methods and human-centered perspectives within the ambit of multimodal pain behaviours recognition, employing an array of deep learning architectures including convolutional neural networks (CNN), long short-term memory networks (LSTM), CNN-Attention networks and CNN with Multi-head Self Attention. The cornerstone of our exploration was to enhance the precision and utility of complex pain recognition endeavors through the lens of statistical relevance and human-centered modality segmentation.
@@ VI Conclusion
Furthermore, the research integrates statistical significance with a human-centered perspective, paving the way for explainable AI in pain recognition. This approach delves into the distinct impact of each data modality on model outcomes, promoting a model of explainability that not only clarifies the workings of complex models but also enhances their adaptability and accuracy. This shift highlights our methodology’s broad applicability, not only in pain management but across diverse domains of human-centered computing, spotlighting its potential to revolutionize how we interact with and leverage AI technologies.
@@
This research intersects affective computing, multimodal data analysis, and pain recognition, aiming to enhance the precision and personalization of pain management through statistical methods. We conscientiously address the potential societal, environmental, and ethical impacts, focusing on aspects such as explainability, transparency, liability, fairness, efficacy, robustness, privacy, security, and sustainability.
@@ Acknowledgment
We’d like to thank Prof. Nadia Berthouze and Dr. Chuang Yu for helpful suggestion. This research is fully supporteds by Holistic AI and partially supported by University College London Interaction Center, University College London, UK.